Chuck wanted to see how much money he uses each week on going out to lunch. Money Spent 17 10 19 20 14 Find the median amount of money Chuck spent on lunch this week.
Answer #1

here putting above data in ascending order:

s. no x
1 9
2 10
3 11
4 14
5 17
6 19
7 20

median =(n+1)/2 th value =(7+1)/2 th value =4th value =14